OpenStudy (ashley1nonly):

The Kp for the reaction below is 1.49 × 108 at 100.0°C: CO(g) + Cl2(g) → COCl2(g) In an equilibrium mixture of the three gases, PCO = PCl2 = 2.22 × 10-4 atm. The partial pressure of the product, phosgene (COCl2), is ________ atm. A) 7.34 B) 3.02 × 1015 C) 3.31 × 10-16 D) 3.31 × 104 E) 6.67 × 1011 Answer: A

OpenStudy (ashley1nonly):

i dont know what i am looking for here

OpenStudy (anonymous):

kk 1 sec

OpenStudy (anonymous):

kk it was the same way we did the 1st problem but this time it is kp, so kp=(COCl2)/(CO)(Cl2)

OpenStudy (ashley1nonly):

and then we solve for COCl2 1.48*10^=x/(2.24*10-4)^2

OpenStudy (anonymous):

Perfect!

OpenStudy (ashley1nonly):

ok let me get one more question

OpenStudy (anonymous):

kk